OpenStudy (stormswan):

could you guys help me please? i just transfered from a different school, and they are learning stuff completely different than in my school. please help. Find -a^2 - 3b^3 + c^2 + 2b^3 - c^2 if a = 3, b = 2, and c = -3.

OpenStudy (anonymous):

notice that -3b^3 + 2b^3 = -b^3 and that c^2 - c^2 = 0 so it is now -a^2 - b^3 simply plug in the values: -(3)^2 - (2)^3 = -9 - 8 = -17

OpenStudy (stormswan):

oh. thank you :) its not like we learn this stuff in Russia...

OpenStudy (anonymous):

glad i could help :)